What is the future value of a 15-year ordinary annuity with annual payments of $2,000 if the annual interest rate is 14%? (Choose the closest value)
Answer #1

[(1+r) -1 FV of Annuity = P P= Periodic Payment r=rate per period n = number of periods

FV = 2000 * (1 + 0.14) 15-1 0.14

FV = 2000 * 43.8424

FV = $87,684.83